Which line is parallel to the line 4x- 5y=10?
a) y=2x-4
b) y=4/5x+8
c) y=-5/4-2
d) y=-2x+10

Found 2 solutions by flame8855, lawdej:
Answer by flame8855(424)

4x- 5y=10
5y = 4x-10
y = (4/5)x -2
the line that is parallel has to have the a slope that is a multiple of 4/5
so b) y=4/5x+8 is the answer

Answer by lawdej(58)

here is the answer
in y=mx+b m is slope
so two say two lines parallel their slope should be the same
rearrange the given equation in the form of y=mx+b
4x- 5y=10
-5y=-4x+10
y=4/5x-2
slope m=4/5
so the answer is choice b